OpenStudy (anonymous):

Write the equation y - 4 = -2(x - 5) in slope-intercept form.

OpenStudy (anonymous):

Do you know what Slope-Intercept form looks like? We use Slope-Intercept Form to show a linear line through an equation that makes the "Slope" and the "Y-Intercept" very easy to see. The form: y = mx+b is called the Slope-Intercept Form because the "m" is the slope, and the "b" is the y-intercept, and very easy too see! So you need to put your equation [which right now is in (Point-Slope) form], into y = mx+b form. Are you with me so far?

OpenStudy (anonymous):

yes, I don't know how to convert.

OpenStudy (anonymous):

Well let's do this, let's isolate y onto the left side of the equation by itself so at least that part looks like y = mx+b, do you know how to do that?

OpenStudy (anonymous):

Add 4 to both sides?

OpenStudy (anonymous):

Exactly!

OpenStudy (anonymous):